Suppose you have a circular coil with a radius of 7 cm that has 50 loops. The way it is oriented makes it form magnetic field lines normal to the area of the coil. If the magnetic field changed from 0.10 T to 0.75 T in 3 milliseconds, what is the average EMF induced in the coil?


1
Expert's answer
2022-04-11T12:07:19-0400
"V=\frac{\Delta \Phi}{\Delta t}=\frac{NA\Delta B}{\Delta t}"

"V=\frac{50*3.14*0.07^2*(0.75-0.10)}{0.003}=167\:\rm V"